OpenStudy (anonymous):

Find the 40th term of the sequence 2, -12, 72, ...

rishavraj (rishavraj):

apply 40th term = a*[r^(n-1)]....n = 40

OpenStudy (anonymous):

an = a1 * r^(n-1) n = number of term to find = 40 a1 = first term = 2 r = common ratio = -6 now we sub a40 = 2 * -6 ^ (40 - 1) a40 = 2 * (-6 ^ 39) a40 = 2 * -2.23 a40 = - 4.46